Let D representing the number of defective units when 5 units are chosen at random from a batch of ten flash drives which contains six defective items. Find the value of random variable D.



1
Expert's answer
2021-01-20T17:45:29-0500

Let "D" be a random var iable defined as ,

"D(x)=" Number of defective flash drives

Then "D" can take the values "0,1,2,3,4,5" .

Now corresponding to each values of "D" we get their probability such as,

"P(D=0)=0"

"P(D=1)=\frac{^6C_1×^4C_4}{^{10}C_5}=\frac{6×1}{252}=\frac{1}{42}"

"P(D=2)=\frac{^6C_2×^4C_3}{^{10}C_5}=\frac{60}{252}=\frac{5}{21}"

"P(D=3)=\frac{^6C_3×^4C_2}{^{10}C_5}=\frac{120}{252}=\frac{10}{21}"

"P(D=4)=\frac{^6C_4×^4C_1}{^{10}C_5}=\frac{60}{252}=\frac{5}{21}"

"P(D=5)=\frac{^6C_5}{^{10}C_5}=\frac{6}{252}=\frac{1}{42}"


Which is required value of random variable "D" .